OpenStudy (anonymous):

Determine if 7, 8, and 13 can be the lengths of the sides of a triangle. If yes, classify the triangle. A. Yes, it is an acute triangle. B. Yes, it is an obtuse triangle. C. Yes, it is a right triangle. D. No, it cannot be a triangle.

ganeshie8 (ganeshie8):

\(\large c^2 = a^2 + b^2 - 2ab \cos (C)\) that gives : \(c^2\gt a^2 +b^2 \implies \text{obtuse}\) \(c^2\lt a^2 +b^2 \implies \text{acute}\) \(c^2= a^2 +b^2 \implies \text{right}\)

ganeshie8 (ganeshie8):

Is \(13^2 > 7^2 + 8^2\) ?

OpenStudy (anonymous):

yes 169 is greater than 113

OpenStudy (anonymous):

so is it B?

ganeshie8 (ganeshie8):

thats right ! but how are you sure that the given lengths indeed form a triangle ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

they could be so small that you wont be able to join the other two sides, to close the shape, right ?

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

you need to check if the given lengths satisfy ALL below conditions : \(a+b < c\) \(b+c < a\) \(c+a< b\)

ganeshie8 (ganeshie8):

In words : `sum of any TWO sides must be greater than the third side`

OpenStudy (anonymous):

Oh yeah i remember that

ganeshie8 (ganeshie8):

good :) B is correct btw ! good job !!

OpenStudy (anonymous):

@ganeshie8 Thank You